OpenStudy (anonymous):

log6(x+5)=12 I think it's not possible because x+5 would have to be negative but not sure

OpenStudy (anonymous):

\[\log_{6}(x+5)=\frac{ 1 }{ 2 } \]

OpenStudy (mathstudent55):

By definitiion of log: 6^(1/2) = x + 5 x = 6^(1/2) - 5 = sqrt(6) - 5

OpenStudy (anonymous):

but sqrt(6)-5 is a negative number. how to find logbase6 of a negative number?

OpenStudy (anonymous):

@mathstudent55

OpenStudy (anonymous):

The condition is x + 5 > 0, not x >0

OpenStudy (anonymous):

I see. I was seeing things incorrectly. thanks